b-) Find the point closest to the origin on the intersection curve of the 2² = 3 (x² + y2) cone and the 2y+42=5 Plane? 2.)
Answer #1

b) Cone! z2 = 3 (x² + y2) Plane! 2y + 4z = 5. on cone Distance from so, we want of any point P(2718) origin rx²+4² + z2 minim& Vi < solving 34 ² = z2 5 Yzt 2y = 5 3-01 y €4-13 +2) = 5 = 3 I 5ß (2 + 4√3) Points A 5 2 (2+48 -5 B co (2 +483) (2-43) (2-4